# Private Federated Learning using Preference-Optimized Synthetic Data

### v1

@@ E.1 Cosine similarity vs Nearest neighbors histogram
E.2 Client selection strategies
Figure 7 shows the effect of changing the number of clients per round and the number of total clients when $\epsilon=7.0$. We find that POPri outperforms the baseline across the board. We also find that POPri is significantly less sensitive to the fraction of clients used per round than the baseline methods. This makes POPri especially useful in settings in which baseline performance may suffer, such as very high client participation regimes.
[…]
E.3 Temperature
Temperature is a key parameter for controlling the diversity of LLM-generated outputs. Increasing the temperature encourages the model to produce less frequent tokens, enhancing diversity. Here we explore the effects of changing temperature in the POPri process and list the results in Table 4. Low temperatures leads to clustering of the text embedding in some regions which does not represent the overall data distribution. However, setting the temperature too high can lead to overly random and potentially incoherent results. Therefore they both lead to lower model accuracy. We therefore choose temperature = 1.0 as our default setting.
Table 4: Ablation results including varying which client ranked-data is chosen as the ‘rejected’ sample for DPO, varying the temperature for the text synthesis process, and varying the preference optimization method.
Rank Temperature Alignment Algorithm
[…]
5th 68.6 0.5 68.1 DPO 68.6
7th 67.7 1.0 68.6 IPO 67.1
10th 66.4 2.0 68.4
E.4 Alignment methods
We also explore IPO (Identity Preference Optimization) (Gheshlaghi Azar et al., 2024), another preference optimization method, to see if it improves algorithm performance. This may help training by alleviating over-fitting, which is a common problem for the DPO algorithm and which affects POPri as well. We show the comparison of next-token prediction accuracy reported by running DPO and IPO algorithm in Table 4. In our case, IPO does not seem to address the overfitting issues and results in worse performance. We therefore choose DPO as our alignment method for finetuning the LLM.
E.5 Rejected sample selection
We construct the DPO preference data via client feedback by generating ten samples from the same prompt and then picking the “selected” and the “rejected” samples. The samples with the highest scores among the ten examples are picked as the “selected” sample in the DPO preference dataset. We experiment on which rank should be utilized as the “rejected” sample in the DPO preference dataset. In Table 4 we show the results of varying which rank is selected for the “rejected” sample. Perhaps surprisingly we find that the 10th rank is not favored. In Fig 8 we further explore the effects by examining the “rejected” and “selected” sample FID scores as a function of round. In the left panel where the “selected” sample FID values are shown, their magnitude and trends behave similarly before they reach the best results (marked by colored dashed vertical lines). For the “rejected” sample FID shown in the right panel, the 5th rank “rejected” samples yield the lowest FID score and therefore smaller gap between the preference sample pairs. However, we also find that higher rank does not always yield better results. This may result from the boundary between the “rejected” and “selected” samples becoming undistinguishable for rank $<$ 5th due to DP noise. We therefore select 5th rank samples as our “rejected” DPO preference samples.
@@ Appendix F Datasets
Dataset ##\## Train Samples ##\## Validation Samples ##\## Test Samples Max Sequence Length Average ##\## of samples per client
@@
We are committed to update the dataset periodically with the latest data to allow future researchers to test their algorithms or ideas against an uncontaminated dataset.

### v2

@@ E.1 Cosine similarity vs Nearest neighbors histogram
Figure 7: In this experiment, we investigate whether the use of label noise-resistant alignment methods could allow the use of higher-ranked rejected samples. To do this, we used the third-ranked sample as the rejected sample, and evaluated different settings for conservative DPO (cDPO) (Mitchell, 2023). We used the bioRxiv dataset experiment setting, set eps=7, learning rate = 8e-7. We find that by tuning the level of conservative-ness we may be able to improve slightly on vanilla DPO.
E.2 Alignment methods
We also experiment with a noise-resistant (or robust) DPO method, conservative DPO (cDPO) (Mitchell, 2023), to see if by using it we can select a higher ranked rejected sample (recall we use the 5th ranked, and higher ranked samples would introduce more noise into the preference pairs). In Figure 7 we find that it can help slightly when choosing a higher rejected sample ranking.
E.3 Rejected sample selection
We construct the DPO preference data via client feedback by generating ten samples from the same prompt and then picking the “selected” and the “rejected” samples. The samples with the highest scores among the ten examples are picked as the “selected” sample in the DPO preference dataset. We experiment on which rank should be utilized as the “rejected” sample in the DPO preference dataset. In Fig 8 we further explore the effects by examining the “rejected” and “selected” sample FID scores as a function of round. In the left panel where the “selected” sample FID values are shown, their magnitude and trends behave similarly before they reach the best results (marked by colored dashed vertical lines). For the “rejected” sample FID shown in the right panel, the 5th rank “rejected” samples yield the lowest FID score and therefore smaller gap between the preference sample pairs. However, we also find that higher rank does not always yield better results. This may result from the boundary between the “rejected” and “selected” samples becoming undistinguishable for rank << 5th due to DP noise. We therefore select 5th rank samples as our “rejected” DPO preference samples.
@@ Appendix F Datasets
Dataset #\# Train Samples #\# Validation Samples #\# Test Samples Max Sequence Length Average #\# of samples per client
@@
We will update the dataset periodically with the latest data to allow future researchers to test their algorithms or ideas against an uncontaminated dataset.